2015 3/A/1/6 loc bridge 97 heard movement approx 50 meters West of bridge, fired illum and spotted one indiv. Engaged with SA fire. Enemy returned fire. Estimated 1 sqd. Fired Arty to West as blocking fire. 2055, Still receiving fire, estimate Plt. Received automatic wpns fire. 1/6 had 1 freindly KIA and got 1 VC KIA(P).
Summary 1/14 continued operations in assigned AO. Alpha and Charlie companies continued S&D in Western and Southern parts of AO. Bravo, Delta and Fox rests and refits in preperation for future operations.
2 November 1967
0210 1/D/1/14 sprung ambush vic 895290. Results: 1 VC KIA. VC was moving to the North.
0725 1/D/1/14 element with an MI team entered a hamlet at 8952282. 3 VC fled from the village. They were engaged with SA fire resulting in 2 VC KIA, 1 VC WIA - CIA. Presently being questioned by MI team. 0800; 1/D/1/14 found 1 VC hiding in bunker, refused to come out, set off Claymore. Results: 1 VC KIA Bunker had 4' entrance, 5' long, 3' wide. VC was dressed in black PJ's, 22 years old, no ID. MI team questioned CIA VC, say he is NVA. He had a Thompson MG which he threw in water. Checking area at this time. The KIA was a guerilla squad leader. Documents on his posession indicated he was a member of the 130th Recon company. The document was dated 29 Oct 67 and also ggave him permission to transport rice. He also had a picture of his unit. CIA led D/1/14 to vic 882282 where they located his M-1 Carbine with 1 magazine.
0945 MSF Med-Caps sighted 2 VC at vic 917254. Pursued to the North. 0955 made contact with 3 VC resulting in 1 VC KIA. Searched hooch and had blocking force move along the highway. 1035 vic 922254 CIA one pack, hammock, medical supplies, bandages and canteen.
1035 1/D/1/14 captured 2 VC in a tunnel, 3 by 10. Searched tunnel vic 894284 found one US frag-type grenade.
Summary 1/14 continued operations in AO Dragon. Coordination for relief of AO Dragon by 1/6 complete. A,B,C/1/6 have relieved elements of 1/14. 1/14 prepares for operations in AO Catmit.
3 November 1967
Summary Golden Dragons continued operations in AO Dragon and prepared to conduct operations in AO Cat Mit.
5 November 67
1320 3/A/1/14 vic 792284 spotted 1 VC walking along a trail going East to West resulting in 1 VC KIA and 1 AK47. VC was wearing khaki uniform and had one poncho, one canteen, w/cover (home made cover), wash cloth, bandages and one hammock.
Summary Golden Dragons continued operations in AO Cat Mit. Charlie lifted from AO Dragon to LZ Storm and commenced operations in present AO after being delayed due to weather. Construction continued on the Firebase at LZ Tempest. Plans to continue firebase construction and conduct S&D operations in AO Cat Mit will continue.
6 November 67
0815 1/D/1/14 vic 754258 spotted one enemy in a bunker. The enemy tried to evade. Was engaged with SA fire resulting in 1 VC KIA, 1 carbine CIA. Enemy was near a bunker. Approx 20-25 bunkers in area large enough to accomodate a company. Searched bunkers and found 1 Chi Com grenade, destroyed in place. Majority of bunkers showed no sign of recent use. KIA was dressed in gray trousers, green shirt. Age approx 25. Kill verified by LT Downs.
1145 2/D/1/14 vic 747255 spotted 3 enemy. Engaged with SA resulting in 1 VC KIA, and 1 M-1 Carbine CIA. They found blood trails and 3 dead VC about 3 to 5 days old. The KIA was wearing black pants and shirt.
1950 1/D/1/14 vic 743258 was probed by an estimated 1 VC. VC set off trip flare. Fired SA and think they hit him. Will check area in morning.
Summary 1/14 continued aggressive S&D operations in the Cat Mit AO. Delta's find confirms intelligence reports of large enemy units infiltrating through the area. Construction of Firebase at LZ Tempest continued. S&D operations will continue in AO Cat Mit.
7 November 67
1227 2/C/1/14 vic 682290 spotted 3 enemy, engaged with SA, resulting in 2 VC KIA and 1 SKS. KIA's wore blue khaki uniforms, ages 25 and 30.
Summary The 1/14 continued S&D operations in AO Cat Mit. A and D are searching what we believe to be a large unit infiltration route. C thinks he is in an area of a seldom but still used infiltration route. 1/14 will continue S&D operations in these areas.
8 November 67
1145 2/A/1/14 vic 786266 found 5 hooches. While searching area they saw some enemy running took them under fire resulting in 3 VC KIA. KIA's were wearing black PJ's. Kill verified by LT Locke.
Summary 1/14 continued S&D operations in AO Cat Mit. B conducts local security sweeps around LZ Tempest. A added 3 enemy KIA's to their score. C and D continued to explore the infiltration routes, selecting night ambush locations. C plans to cover as much ground as possible. A and D plan to use daylight ambushes.
9 November 67
0845 1/A/1/14 vic 803268 while searching area spotted 1 enemy. Took him under fire resulting in 1 VC KIA, wearing black PJ's, and carrying a pack with cooking utensils and rice.
1100 3/4 LRRP vic 704282 spotted 1 VC, engaged enemy and saw him fall. At the same time 2 more VC ran out of bushes. Called in Arty and checked area.
Summary Golden Dragons continued S&D operations in AO Cat Mit. C combat assaulted to the North East in order to work more terrain in less time. A and D will continue to work the thick terrain, covering the valley draws and hilltops. B minus will continue security of LZ Tempest, with one platoon engaged in S&D operations.
10 November 1976
0945 1/D/1/14 vic 725273 spotted 4 enemy, took them under fire, resulting in 4 VC KIA, dressed in black PJ's.
1045 2/A/1/14 vic 785264 spotted 1 enemy and set up an ambush resulting in 1 VC KIA and 1 Russian Carbine.
Summary 1/14 continues S&D operations in AO Cat Mit. A is responsible for another find that indicates the presence of a large enemy element. A and D will continue operations in that area, attempting to make contact. C will continue S&D operations in their AO. B continues it's combined security and S&D missions.
11 November 1967
Summary 1/14 continued S&D operations in AO Cat Mit. A, C, and D continued S&D in areas of previous contact, conducting daylight and night ambushes.
12 November 1967
0940 2/A/1/14 made contact vic 785265. 1 friendly WIA. 1/A.1.14 moved to help 2/A/1/14. US WIA hit in side. Think 1 man did the shooting. Found 2 M-16 magazines.
1045 D/1/14 vic 755255 killed 1 VC with pack. Dressed in black PJ's, approx 20 years old.
1345 3/C/1/14 vic 688319 spotted 5 to 6 enemy moving north. 3/C/1/14 set up ambush. Engaged enemy with SA, fired Arty to block escape. Expended Gunships on enemy position. 1 enemy KIA, 1 AK47 CIA. C.O. 1/14 lifted 2/C.1.14 to South of contact area and they killed 2 enemy. Total results: 3 enemy KIA dressed in green uniforms, 3 packs and pith helmets.
1447 D/1/14 vic 750255 spotted approx 15 enemy. Engaged with SA and fired Arty. Engaged with Gunships. Swept area and found 5 VC KIA, 2 NVA CIA. 3 KIA wearing green uniforms and pith helmets. 2 wearing mixed uniforms. Kill verified by Cpt Sells.
1705 2/C/1/14 vic 688319 spotted enemy. Engaged with SA fire resulting in 1 enemy KIA. Kills verified by LT Barry and LT Ensign.
Summary 1/14 continued S&D operations South-West of Duc Pho district in AO Cat Mit. D's contact with a large party carrying mortar ammo is another indication of a large force being in the area. A, C, and D continued S&D operations in AO. B continued FSB security and local S&D operations.
13 November 67
1055 1/D/1/14 vic 768249 received SA fire from 1 VC. 1 friendly WIA (arm). Returned fire with SA and M-79. Sighted 3 VC, engaged with SA. Results: 2 VC KIA.
1400 3/D/1/14 vic 766247 has 1 enemy KIA. KIA approx 20-22 years old.
1450 1/D/1/14 vic 770248 has 1 enemy KIA wearing green uniform.
1655 1/D/1/14 has 1 enemy KIA vic 772251 wearing green uniform.
Summary 1/14 continued their aggressive and effective S&D operations in AO Cat Mit. A and D continued to work in S&D portion of AO. C combat assualted to the Northwest in an attemp to gain further contact. B continued local security patrols vic LZ Tempest.
14 November 1967
Summary 1/14 continued S&D operations in AO Cat Mit. A and D continued S&D in South and Southeast portions of AO respectively. C worked in Northeast sections. All conducted daylight patrols and ambushes. B continued FSB security, Bn continued aggressive operations in assigned AO.
15 November 1967
1142 2/C/1/14 contact vic 682288. Results: 3 NVA KIA, 18-23 years of age, wearing grey khaki uniforms.
1305 2/A/1/14 vic 785246 spotted 2-3 enemy and pursued. 2 enemy KIA, 2 wpns CIA. Enemy wearing green uniforms and black PJ's.
Summary The Golden Dragons continued S&D operations in assigned AO. The contact made and type of equipment seem to indicate further presence of a large element. A, C, and D continued to S&D in their assigned AO. B continued combined FSB security and S&D operations.
16 November 1967
Summary Golden Dragons continued S&D operations in AO CAT MIT. LZ Tempest received mortar attack. Plans are to continue S&D operations in southern portion of AO CAT MIT.
17 November 1967
1144 1/A/1/14 vic 785244 had 4 US WIA. 3 walking and 1 litter.
Summary The Golden Dragons continued operations in AO CAT MIT. Charlie Co was extracted and occupied positions in Southern portion of AO DRAGON. LZ Tempest received 14-17 rds incoming 82mm rds. Negative damage was reported as most of rds fell outside of the perimeter. One NVA grenadier was killed by his own mortar rds at the wire with 7 grenades. Plans to continue S&D operations in area of airstrikes.
18 November 1967
1325 3/A/1/14 vic 784234 while sweeping area, spotted movement. Engaged with SA fire and received return fire. Called in Arty blocking fire. Results: 1 enemy KIA wearing green uniform.
Summary The GOLDEN DRAGONS continued operations from LZ Tempest to destroy enemy forces. Alpha and Delta conducted joint operations to determine the location of NVA base area. Air and Artillery was used against a stubborn enemy. Plans to continue S&D operations with Alpha and Delta continue.
19 November 1967
1345 1/D/1/14 in contact with 5-6 enemy vic 789233. Arty blocking fire called in. 1 82mm mortar CIA. 1410; 1 more 82mm mortar CIA. 1415; 3 enemy KIA.
1613 1/D/1/14 vic 790231 has 1 enemy KIA. 2/D/1/14 vic 785231 has 1 enemy KIA, both wearing khaki uniforms.
1712 3/A/1/14 vic 790231 hit a landmine. 4 US WIA, 1 head injury, 1 foot blown off, 1 not too serious, 1 leg injury.
Summary GOLDEN DRAGONS continued S&D operations with 2 rifle companies for the base area of the 2nd NVA Reg. D Co found a hooch complex and captured 2 complete 82mm mortars and killed 5 NVA and captured 6 weapons. A Co suffered 4 US WIA from booby traps while departing the same hooch complex. Plans for the exchange of Alpha and Charlie are in preparation. S&D operations will continue with 2 rifle companies.
20 November 67
0725 1/A/1/14 contact at 785241. 1 enemy KIA, 1 AK47 CIA. Indiv wearing NVA uniform, web gear, 2 chicom grenades.
0855 3/D/1/14 vic 790229 has 1 VC KIA, 1 AK-47 CIA. Indiv dressed in black PJ's and blue sweater. Had 1 US frag grenade and 3 AK-47 magazines.
1925 3/B/1/14 vic 757260 had 1 enemy walk into their ambush position. Results: 1 enemy KIA. Indiv wearing black PJ's. Will check area in morning.
Summary GOLDEN DRAGONS continued operations in AO CAT MIT with 3 rifle companies. C Co was extracted from Sa Huyn and replaced by A Co. Charlie and Delta will continue S&D operations in the Southern portion of AO CAT MIT.
21 November 67
1200 3/D/1/14 vic 775224 has 1 NVA KIA. Indiv wearing green shorts and shirt. Also pursued 1 other NVA. 1330; swept area with neg findings.
1635 2/C/1/14 vic 777236 found 1 enemy KIA, 1-2 weeks old. Nude with a khaki uniform to one side of body. Killed by Arty.
Summary GOLDEN DRAGONS continued operations (S&D) in assigned AO. Co A continued OPCON to 2/35 in AO DRAGON. B Co provided security for LZ Tempest. C Co and D Co to continue S&D operations.
22 November 67
1050 1/D/1/14 vic 781227 sustained 1 US KIA and 1 US WIA. Believe enemy is dug in with a machine gun.
1338 1/D/1/14 has 1 enemy KIA, 1 AK-47 CIA vic 781227.
1342 2/D/1/14 has 1 enemy KIA wearing NVA uniform.
1400 3/D/1/14 vic 783229 has 1 enemy (NVA) CIA wearing NVA uniform and blue sweatshirt
1411 1/B/1/14 vic 737286 found built up area with 11 hooches, VN newspaper dated June '67. Also 1 enemy killed by Arty H&I fire, (B/2/9) approx 1 week old. Enemy dressed in black PJ's with military haircut. Area not used by any significant sized force in past several months.
1520 C/1/14 vic 782242 spotted approx 30 enemy moving toward their location. C/1/14 set up an ambush, gunships requested, Arty laid on.
1642 C/1/14 still in contact has 5 enemy KIA.
1650 C/1/14 KIA's as follows: 1600, CP-1 NVA KIA; 1700 CP - 1 NVA KIA; 1745, 3rd plt - 2 NVA KIA; 1900, 3rd plt - NVA KIA; enemy KIA total for 22 NOV: 10.
Summary Operations continued in AO CAT MIT by the GOLDEN DRAGONS. C Co made contacts with the 132nd Signal Company and captured a Chicom radio receiver. D Co found and destroyed a signal units complex of hooches and captured a field telephone, batteries and radio tubes. B Co continued security of LZ Tempest. Plans: C&D companies continued S&D operations in vicinity of area of contact.
23 November 67
0725 C/1/14 vic 785420 found 1 indiv hiding in bushes. Pack and web gear, dressed in blue khaki pants.
1200 3/C/1/14 vic 780239 have 1 enemy KIA. KIA had 1 pack wearing grey khaki uniform.
1435 D/1/14 had 1 man slightly wounded by an M-79 cannister round. Man was moving down a trail when the M-79 accidentally went off and wounded him in the foot.
Summary GOLDEN DRAGONS continued operations in AO CAT MIT with 3 rifle companies. C and D companies continued with units receiving turkey dinners. Plans: Continuation of S&D operations in suspected 2nd NVA base area.
24 November 67
0734 3/C/1/14 vic 786240 found 1 shallow grave of an NVA. Dead approx 2-3 days. Indiv killed by small arms.
0950 2/C.1.14 vic 790232 made contact with approx 4-5 enemy. Results: 2 NVA KIA, dressed in green uniforms, 2 packs. Swept area, called for Arty blocking fire.
1050 3/C/1/14 vic 793237 had 1 enemy KIA and 1 AK-47 CIA.
1220 2/C/1/14 vic 789230 found 2 enemy KIA by Arty (B/2/9), wearing grey-green khaki uniforms. Neg weapons or web gear. Dead approx 3 hours.
1425 1/C/1/14 vic 798237 spotted 2 enemy moving through brush. Engaged with SA fire and swept area with neg results.
Summary Operations continued in AO CAT MIT with 3 rifle companies. S&D operations continued with C and D companies in Southeastern portion of AO. Plans: Continue operations in the suspected 2nd Reg base area.
25 November 67
0930 C/1/14 vic 800233 found a fresh grave with 1 enemy dead as a result of SA. Shot in abdomen sometime 24 NOV 67.
1429 3/D/1/14 vic 783208 has 1 enemy KIA, 1 SKS CIA, one pack, wearing green NVA uniform and Ho Chi Minh sandals.
Summary Operations continued in AO CAT MIT with 2 rifle companies. C Co made contact with 1 NVA and captured 1 SKS after receiving AW fire from 3 locations. Plans: Continue S&D operations in the same general area.
26 November 1967
1046 2/D/1/14 vic 803213 spotted 4 enemy at a distance of 350 meters. Engaged with SA. Results: 2 NVA KIA, 2 packs CIA. Wearing green uniforms, neg wpns. 1250 found 1 enemy KIA wearing green pants, black shirt.
1410 1/C/1/14 vic 795220 had 2 NVA KIA and 1 AK47. More enemy in open. Called in Arty.
1710 At 1345 vic 795227 1/C/1/14 observed 2 enemy moving. Engaged with SA. 2 NVA KIA wearing green khaki uniforms and blue shirts. KIA's approx 18-26 yrs old.
1711 2/C/1/14 vic 803229 heard noises in brush. fired SA. Searched area for body and found 1 enemy KIA, dressed in green uniform and blue shirt.
Summary GOLDEN DRAGONS continued operations in AO CAT MIT. D Co was extracted to provide security for LZ Tempest. B Co was airlifted to the AO and will conduct S&D operations to the west. C Co continued S&D operations in the Southern portion of the AO. Plans: Continue S&D operations in AO CAT MIT with 3 companies.
27 November 67
Summary GOLDEN DRAGONS continued operations in AO Cat Mit with 3 rifle companies. Contact was light with C Company making light contact. GOLDEN DRAGONS had negative enemy kills or US casualties. Plans: Continue S&D operations with 2 rifle companies and fire base security with 1 rifle company.
28 November 67
1315 2/B/1/14 vic 786203 has 1 enemy KIA, wearing green fatigues.
Summary 1st 14th Infantry continued S&D opns in AO Cat Mit with 3 rifle companies and one reconnaissance platoon. Continued heavy rain and overcast hampered operations as enemy movement was limited. Weather prevented the arrival of the C&C aircraft throughout the day. Sight contact was made by B Company. An enemy radio transmission was again reported in the AO. Plans: Continuation of S&D opns with Bravo and Charlie companies. 1 platoon of Delta will conduct CA with Hoi Chanh to weapons cache.
29 November 67
1003 Bravo 20 has one NVA KIA.
1110 Charlie 36 at BS 785241 saw 2 NVAs coming up a trail from the Northwest. They set up a custer and one man took both NVA under fire. Two NVA KIA.
1327 Bravo 30 at 1030 hrs spotted three NVA. Took them under fire at BS 781214. Has oone NVA KIA and one SKS CIA.
Summary GOLDEN DRAGONS continued S&D opns in the high ground in the southern portion of AO Cat Mit with 2 rifle companies and a reconnaissance platoon. Charlie Company swept toward the area of the Spot Report. Delta conducted a platoon sized CA to the area of the weapons cache accompanied by Hoi Chanh. Results where negative. Charlie company located a training area complex with 30 bunkers underneath the hooches. Plans: Continue operations to destroy the 22d NVA Regimental CP.
30 November 1967
1000 Charlie 16 while sweeping the contact area of yesterday found 1 fresh shallow grave. The 16 dug it up and found 1 NVA dressed in light green uniform, killed by gunships.
Summary GOLDEN DRAGONS continued operations in AO Cat Mit with 2 rifle companies. Weather completely closed into LZ Tempest and continued during the period with heavy rains. No contact was reported. Plans: Alpha Company will be picked up at LZ Charlie Brown and Bridges 99 & 95 and inserted into AO Cat Mit. Charlie Company will be extracted and will return to LZ Charlie Brown. Bravo company and Alpha will continue S&D opns to the west in the AO.
There are no Journals for December, 1967 in the Archives